OpenStudy (anonymous):

Identify the Newton's law that best explains this situation- A hand pases through a piece of wood in a karate chop. (Please explain your reasoning)

OpenStudy (anonymous):

newton's 2nd law. F=(p2-p1)/(t2-t1) or F(t2-t1)=p2-p1=constant if t2-t1 is small then F is large......

OpenStudy (anonymous):

can you please explain this.

OpenStudy (anonymous):

this is also known as impulse. Ft=rate of change of momentum=constant since the product of F and t is constant so as F increases t decreases and vice-versa.